OpenStudy (anonymous):

Math Help Please!!!! Does the data in the table represent a direct variation or an inverse variation? Write an equation to model the data in the table. X | 2 | 4 | 8 | 12 Y | 6 | 3 | 3/2 | 1 A) inverse variation; xy = 12 B) direct variation; y = 12x C) direct variation; y = 12/x D) inverse variation; y/x = 12

OpenStudy (anonymous):

@Daniellelovee @Godlovesme

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (daniellelovee):

which one do you think it is?

OpenStudy (anonymous):

I don't have any clue that's why I need some help on it.

OpenStudy (daniellelovee):

well notice how all of the options = (now tell me what do they equal)

OpenStudy (anonymous):

12 in some way

OpenStudy (daniellelovee):

exactly and they are being multiplied therefore the answer is....

OpenStudy (anonymous):

12.....

OpenStudy (daniellelovee):

no lol I meant A,B,C, or D

OpenStudy (anonymous):

xy = 12 and the xy showing that they are being multiplied.. but how do you know its inverse variation??

OpenStudy (xapproachesinfinity):

seems to me an inverse variation inverse variations are of the form \[y=\frac{k}{x}\] here you have \[y=\frac{12}{y}\]

OpenStudy (xapproachesinfinity):

i mean over x not over y lol

OpenStudy (anonymous):

So the answer is D and not A ?

OpenStudy (xapproachesinfinity):

your answers somewhat are not constructed nicely well xy=12 is inverse variation as well

OpenStudy (xapproachesinfinity):

and why is D called a direct variation something is wrong there

OpenStudy (xapproachesinfinity):

please make sure about the problem and post it neatly !!!

OpenStudy (anonymous):

D and A say inverse variation !!

OpenStudy (daniellelovee):

D sais inverse but the answer is A because it matches the equation

OpenStudy (xapproachesinfinity):

you wrote D direct variation my dear

OpenStudy (anonymous):

Does the data in the table represent a direct variation or an inverse variation? Write an equation to model the data in the table. X | 2 | 4 | 8 | 12 Y | 6 | 3 | 3/2 | 1 A) inverse variation; xy = 12 B) direct variation; y = 12x C) direct variation; y = 12/x D) inverse variation; y/x = 12

OpenStudy (anonymous):

It is the same way it is listed above. Anyways, thanks @Daniellelovee

OpenStudy (xapproachesinfinity):

for me both A) and D are correct

OpenStudy (xapproachesinfinity):

post the snapshot of the problem otherwise your question has errors

OpenStudy (xapproachesinfinity):

here for more insight \[xy=k \Longrightarrow y=\frac{k}{x}, or~~x=\frac{k}{y}\] this show that x can vary inversely with y or y can vary inversely with x